# Do Schwartz Higher-Order Values Help Sentence-Level Human Value Detection? When Hard Gating Hurts

### v1

@@ 5.1 Limitations and threats to validity
Our findings have several limitations. First, we report single-run results, so differences of 1–2 Macro- points may be unstable. Second, the sentence-level setting limits signal; hierarchy may help more with broader discourse context. Annotation noise and multi-label overlap can also conflict with strict parent–child constraints. Third, calibration can overfit under severe imbalance, especially for rare categories (e.g., Openness). Finally, conclusions are scoped to this benchmark and a compute-frugal regime; results may shift with more data, new domains, alternative schemes, or larger models.
@@ 6 Conclusions and future work
This paper examined whether higher-order (HO) value abstractions improve fine-grained, sentence-level human value detection under a compute-frugal regime. Across encoder-based systems and compact instruction-tuned LLMs, we tested HO-aware strategies spanning direct HO prediction, HOvalue inference via hard gating/cascades, and low-cost upgrades such as label-wise threshold calibration and small soft-voting ensembles.
Our results yield three conclusions. First, HO categories are learnable from single sentences, but learnability is uneven and correlates with prevalence and lexical regularities: pairs such as Growth/Self-Protection are easier than rare or diffuse categories such as Openness, which also shows persistent pole asymmetries. Second, the most reliable gains come from calibration and ensembling, not from handcrafted features or strict hierarchical inference. Label-wise threshold tuning helps often (but can overfit under severe imbalance), and small soft-voting ensembles provide the most consistent gains. Third, hard hierarchical mechanisms look strong under conditional evaluation but do not robustly improve the end task. Presence gating and HOvalues masking raise validation scores on gate-passing sentences, but gains largely disappear (or reverse) on the full distribution, consistent with error compounding and recall suppression under sparse supervision. HO structure is informative, but enforcing it as a hard constraint is often too brittle for noisy, multi-valued sentences.
Compact instruction-tuned LLMs are not competitive with supervised encoders in absolute Macro- under the same budget, but they can still help as complementary signals: in some slices, cross-family ensembles benefit from LLM diversity even when LLMs lag as standalone predictors. Overall, hierarchy is best used as an inductive bias that guides inference softly, rather than as a rigid routing rule, at least for sentence-level value detection with sparse multi-label targets.
Future work should replace brittle hard gates with approaches that preserve uncertainty and better match the annotation reality. A promising direction is joint hierarchical learning, where HO and fine-grained labels are predicted in a single model via multi-task objectives or structured decoders with shared representations. Closely related are soft HO priors that condition value predictions on HO probabilities (rather than masking), which can encourage coherence without catastrophic recall loss. Given instability under imbalance, more principled calibration for rare labels is also needed. Finally, because sentence-level inputs limit signal and values may be expressed across discourse, future work should systematically vary context (sentence, local window, document) to test which values benefit from broader evidence. Evaluations on additional domains and schemes will help clarify when value hierarchies help most and how to use them robustly in data-limited settings.
7 CRediT authorship contribution statement
Víctor Yeste: Conceptualization, Methodology, Software, Validation, Formal analysis, Investigation, Resources, Data Curation, Writing - Original Draft, Writing - Review & Editing, Visualization, Project administration. Paolo Rosso: Supervision, Writing - Review & Editing.
@@ 8 Declaration of generative AI and AI-assisted technologies in the manuscript preparation process
During the preparation of this work, the authors used ChatGPT from OpenAI to improve the readability and language of the manuscript. After using this tool/service, the authors reviewed and edited the content as needed. The authors take full responsibility for the content of the published article.
@@ 9 Data availability
This study uses the English, machine-translated ValueEval’24/ValuesML release (Mirzakhmedova et al., 2024; The ValuesML Team, 2024). The dataset is distributed under a Data Usage Agreement that allows research use but does not permit redistribution of the texts. Accordingly, we cannot share the sentences or any derivative file that contains original textual content. Researchers with appropriate access can obtain the official train/validation/test splits by registering and downloading the release from Zenodo (The ValuesML Team, 2024).
@@ Appendix A Value-to-HO mapping
Table 1 reports the fixed mapping used in this paper between Schwartz’s 19 refined basic values and the eight HO categories. Note that in the refined theory some basic values contribute to more than one HO category (e.g., Hedonism, Achievement, Face, Humility), so overlaps across rows are expected.

### v2

@@ 5.1 Limitations and threats to validity
Our findings have several limitations. First, we report single-run results, so differences of 1–2 Macro- points may be unstable. Second, the sentence-level setting limits signal; hierarchy may help more with broader discourse context. Annotation noise and multi-label overlap can also conflict with strict parent–child constraints. Third, calibration can overfit under severe imbalance, especially for rare categories (e.g., Openness). Finally, this is a benchmark-driven study: the conclusions are tied to ValueEval’24 / ValuesML and a compute-frugal regime, so external validity across domains, languages, annotation schemes, or larger-model settings remains to be established in future work.
@@ 6 Conclusions and future work
This paper presented a compute-bounded empirical study of whether higher-order (HO) value abstractions improve sentence-level human value detection. The results show that HO categories are learnable, but their difficulty varies markedly with prevalence and lexical regularity. The most reliable improvements come from label-wise threshold calibration and small soft-voting ensembles, whereas hard hierarchical mechanisms such as Presence gating and HOvalue masking do not robustly improve the end task despite looking stronger under conditional evaluation. Compact instruction-tuned LLMs also remain weaker than supervised encoders in absolute Macro-, although they can still add useful diversity in some cross-family ensembles. Overall, the benchmark-level conclusion is that HO structure is useful as an inductive bias, but too brittle when enforced as a hard routing rule for sparse, noisy, multi-label sentence classification.
Future work should therefore focus on hierarchy-aware methods that preserve uncertainty instead of discarding it. Promising directions include joint hierarchical learning, soft HO priors that condition value predictions on HO probabilities rather than binary masks, and stronger calibration methods for rare labels. Because sentence-level inputs often underrepresent value cues, future studies should also vary the amount of context available to the model and test these approaches across additional domains and annotation schemes. In other words, external validity beyond this benchmark-driven setting remains an open question for future work.
@@ 7 Declaration of generative AI and AI-assisted technologies in the manuscript preparation process
During the preparation of this work, the authors used ChatGPT from OpenAI in order to improve the readability and language of the manuscript. After using this tool/service, the authors reviewed and edited the content as needed and take full responsibility for the content of the published article.
@@ 8 Data availability
This study uses the English, machine-translated ValueEval’24/ValuesML release (Mirzakhmedova2024; ValueEval24Zenodo). The dataset is distributed under a Data Usage Agreement that allows research use but does not permit redistribution of the texts. Accordingly, we cannot share the sentences or any derivative file that contains original textual content. Researchers with appropriate access can obtain the official train/validation/test splits by registering and downloading the release from Zenodo (ValueEval24Zenodo).
@@ Appendix A Value-to-HO mapping
Table A1 reports the fixed mapping used in this paper between Schwartz’s 19 refined basic values and the eight HO categories. Note that in the refined theory some basic values contribute to more than one HO category (e.g., Hedonism, Achievement, Face, Humility), so overlaps across rows are expected.
@@ Appendix B Label prevalence across data splits
Appendix C Prompt templates for reported LLM experiments
This appendix records the prompt format used for the prompted-LLM results reported in Section 4.6. The prompt family was inherited from the controlled prompt comparison in Yeste2026, where the definition-style template outperformed simpler direct, QA-style, and hidden-CoT variants under the same compute-frugal regime. We therefore document the best-performing template used in this paper rather than re-running a full prompt-engineering sweep.
System prompt.
You are a moral-psychology assistant. Using the refined basic values taxonomy (Schwartz 1992; Schwartz et al. 2012), answer the user’s labeling requests exactly as instructed.
Zero-shot user prompt (definition style).
### Value definitions
- <value_1>: <one-line definition>
- <value_m>: <one-line definition>
### Task
Identify which of the above values the SENTENCE relates to. Return only a JSON array of the matching value names.
SENTENCE: <sentence>
The definition block contains one-line descriptions of the candidate labels derived from Schwartz2012overview. For slice-restricted runs, the candidate list is limited to the values relevant to that slice. For example, one Self-Enhancement run used the label subset [Hedonism, Achievement, Power: dominance, Power: resources, Face].
Few-shot wrapper.
SENTENCE: <example sentence>
OUTPUT: <example JSON array>
For few-shot prompting, exemplars with the above schema are prepended before the final zero-shot query, with . When , the exemplar pool includes at least one null example whose output is [].
Decoding and parsing.
All prompted runs use greedy decoding with max_new_tokens=200. Outputs are parsed as JSON arrays and mapped to labels by exact string matching. Invalid JSON outputs or out-of-vocabulary labels are treated as empty predictions. For HO evaluation, predicted basic values are mapped to HO labels via Eq. (1).